# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v1

@@ 3.3. Vision and failure of stronger clone
Such a decomposition is guaranteed to exist for any local randomizer. However, an error occurred in the construction of the reduction pair $P_{0}$ and $P_{1}$ based on this decomposition. Similar to the standard clone in Section 3.1, they define the following distribution $P_{0}(\epsilon_{0},p)$ and $P_{1}(\epsilon_{0},p)$: For any $p\in[0,1/(e^{\epsilon_{0}}+1)]$, let
@@
$\Delta_{1}\sim\text{Bern}(e^{\epsilon_{0}}p),\quad\Delta_{2}\sim\text{Bin}(1-\Delta_{1},p/(1-e^{\epsilon_{0}}p)).$
@@
$P_{0}(\epsilon_{0},p)=(A+\Delta_{1},C-A+\Delta_{2})\quad\text{and}\quad P_{1}(\epsilon_{0},p)=(A+\Delta_{2},C-A+\Delta_{1}).$
@@
$D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\epsilon}}\big{(}P_{0}(\epsilon_{0},p),P_{1}(\epsilon_{0},p)\big{)},$
which serves as an upper bound for a specific local randomizer (different $\epsilon_{0}$-DP randomizers may have different values of $p$). Leveraging Lemma 7, they would then conclude the general upper bound for any $\epsilon_{0}$-DP local randomizer:
(1) $D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\epsilon}}\big{(}P_{0}(\epsilon_{0},\frac{1}{e^{\epsilon_{0}}+1}),P_{1}(\epsilon_{0},\frac{1}{e^{\epsilon_{0}}+1})\big{)}.$
Lemma 7 (Lemma 5.1. in (Feldman et al., 2023a)).
@@
$D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p)\,\|\,P_{1}(\varepsilon_{0},p)\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p^{\prime})\,\|\,P_{1}(\varepsilon_{0},p^{\prime})\big{)}$
Unfortunately, they encountered difficulty in constructing a post-processing function $f$ for this construction of $P_{0}$ and $P_{1}$. While they provided a function in the original paper, it was proven to be incorrect in the corrected revision (Feldman et al., 2023b). The issue arises from the fact that the “leftover” distribution of $x_{1}$ (i.e., $\mathcal{Q}_{1}$) is mixed with the “leftover” distribution of $x_{i}$ (i.e., $\mathcal{Q}_{i}$) in the above construction. In this case, the function $f$ does not know which distribution to sample from. This technical problem is fundamental and remains unsolved.
Although the corrected version was published on arXiv in October 2023, this error has been propagated in subsequent works (Wang et al., 2024; Chen et al., 2024; Yang et al., 2024b; Wang et al., 2025). For instance, the variation-ratio framework made significant efforts to design an algorithm to find the parameter $p$ for various specific local randomizers in the decomposition of Theorem 6. However, their work relies on the incorrect post-processing function $f$ presented in (Feldman et al., 2023a), which renders their results invalid.
Due to this fundamental difficulty, it is required that $x_{1}$ has no “leftover” distribution. In other words, each component of $\mathcal{R}(x_{1})$ must be distinguishable from the “leftover” distribution of $x_{i}$ for $i\geq 2$. In the above example, this necessitates a four-point-based construction for $P_{0}(\epsilon_{0},p,q)$ and $P_{1}(\epsilon_{0},p,q)$ (Feldman et al., 2023b):
$\mathcal{R}(x_{1}^{0})$ $=e^{\epsilon_{0}}p\mathcal{Q}_{1}^{0}+p\mathcal{Q}_{1}^{1}+(1-p-e^{\epsilon_{0}}p)\mathcal{Q}_{1},$
$\mathcal{R}(x_{1}^{1})$ $=p\mathcal{Q}_{1}^{0}+e^{\epsilon_{0}}p\mathcal{Q}_{1}^{1}+(1-p-e^{\epsilon_{0}}p)\mathcal{Q}_{1},$
@@
This new decomposition proposed in the corrected version introduces additional challenges. First, for specific randomizers, computing a tight value of $q$ is nontrivial. Second, the monotonicity of $D_{e^{\epsilon}}\big{(}P_{0}(\epsilon_{0},p,q),P_{1}(\epsilon_{0},p,q)\big{)}$ with respect to $p$ is not known. Consequently, we are unable to derive the desired conclusion—namely, a general upper bound applicable to any $\epsilon_{0}$-DP local randomizer, as stated in Formula (1). For the same reason, it remains unclear whether this decomposition necessarily yields tighter bounds than the standard clone decomposition. More critically, the new bound $D_{e^{\epsilon}}\big{(}P_{0}(\epsilon_{0},p,q),P_{1}(\epsilon_{0},p,q)\big{)}$ for a specific local randomizer lacks an efficient algorithm to compute.
4. General Clone and the Optimal Bounds
@@
The hierarchy of the bounds provided by the decomposition-based methods is shown in Figure 1.
Figure 1. Hierarchy among decompositions-based methods
4.1. Definition of general clone
Definition 0 (Decomposition in the General Clone Paradigm).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a local randomizer. The general clone paradigm considers the following decomposition:
@@ 4.1. Definition of general clone
(2) $\forall i\in[2,n],\quad\mathcal{R}(x_{i})$ $=\sum_{j=1}^{k}c_{j}Q_{1}^{j}+\beta Q_{i},$
where $Q_{1}^{j}$ for $j=1,\dots,k$ and $Q_{i}$ for $i=2,\dots,n$ are probability distributions over $\mathbb{Y}$, and $a_{j},b_{j},c_{j},\beta$ are non-negative coefficients satisfying:
$\sum_{j=1}^{k}a_{j}=1,\quad\sum_{j=1}^{k}b_{j}=1,\quad\sum_{j=1}^{k}c_{j}+\beta=1.$
The general clone paradigm characterizes the general form of decompositions used in privacy amplification analysis. It directly subsumes the decomposition used in the standard clone framework. Although the decomposition defined by the privacy blanket appears structurally different, we will show that it naturally corresponds to a valid decomposition under the general clone paradigm.
@@
$\Pr[A_{0}=j]=\begin{cases}a_{j}&\text{for }j\leq k,\\ 0&\text{for }j=k+1.\end{cases},\Pr[A_{1}=j]=\begin{cases}b_{j}&\text{for }j\leq k,\\ 0&\text{for }j=k+1.\end{cases}$
@@
The reduction pair $(P_{0}^{GC},P_{1}^{GC})$ (with “GC” denoting “General Clone”) is defined as the distributions of the histograms over $\{1,\dots,k+1\}$ generated by sampling:
  • One sample from $A_{0}$ for $P_{0}^{GC}$ (or $A_{1}$ for $P_{1}^{GC}$);
@@
4.2. General clone is not stronger than blanket

### v5

@@ 3.3 Vision and failure of stronger clone
Such a decomposition is guaranteed to exist for any local randomizer. However, an error occurred in the construction of the reduction pair $P_{0}$ and $P_{1}$ based on this decomposition. Similar to the standard clone in Section 3.1, they define the following distribution $P_{0}(\varepsilon_{0},p)$ and $P_{1}(\varepsilon_{0},p)$: For any $p\in[0,1/(e^{\varepsilon_{0}}+1)]$, let
@@
$\Delta_{1}\sim\text{Bern}(e^{\varepsilon_{0}}p),\quad\Delta_{2}\sim\text{Bin}(1-\Delta_{1},p/(1-e^{\varepsilon_{0}}p)).$
@@
$P_{0}(\varepsilon_{0},p)=(A+\Delta_{1},C-A+\Delta_{2})\quad\text{and}\quad P_{1}(\varepsilon_{0},p)=(A+\Delta_{2},C-A+\Delta_{1}).$
@@
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p)\parallel P_{1}(\varepsilon_{0},p)\big{)},$
which serves as an upper bound for a specific local randomizer (different $\varepsilon_{0}$-DP randomizers may have different values of $p$). Leveraging Lemma 2, they would then conclude the general upper bound for any $\varepsilon_{0}$-DP local randomizer:
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},\frac{1}{e^{\varepsilon_{0}}+1}),P_{1}(\varepsilon_{0},\frac{1}{e^{\varepsilon_{0}}+1})\big{)}.$ (1)
Lemma 2 (Lemma 5.1. in [FMT23a]).
@@
$D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p)\parallel P_{1}(\varepsilon_{0},p)\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p^{\prime})\parallel P_{1}(\varepsilon_{0},p^{\prime})\big{)}$
Unfortunately, they encountered difficulty in constructing a post-processing function $f$ for this construction of $P_{0}$ and $P_{1}$. While they provided a function in the original paper, it was proven to be incorrect in the corrected revision [FMT23b]. The issue arises from the fact that the “leftover” distribution of $x_{1}$ (i.e., $\mathcal{Q}_{1}$) is mixed with the “leftover” distribution of $x_{i}$ (i.e., $\mathcal{Q}_{i}$) in the above construction. In this case, the function $f$ does not know which distribution to sample from. This technical problem is fundamental and remains unsolved.
Although the corrected version was published on arXiv in October 2023, this error has been propagated in subsequent works [WPL+24, CCG24, YYF+24, WZL+25, KHH23]. For instance, the variation-ratio framework made significant efforts to design an algorithm to find the parameter $p$ for various specific local randomizers in the decomposition of Theorem 3. However, their work relies on the incorrect post-processing function $f$ presented in [FMT23a], which renders their results invalid.
Due to this fundamental difficulty, it is required that $x_{1}$ has no “leftover” distribution. In other words, each component of $\mathcal{R}(x_{1})$ must be distinguishable from the “leftover” distribution of $x_{i}$ for $i\geq 2$. In the above example, this necessitates a four-point-based construction for $P_{0}(\varepsilon_{0},p,q)$ and $P_{1}(\varepsilon_{0},p,q)$ [FMT23b]:
$\mathcal{R}(x_{1}^{0})$ $=e^{\varepsilon_{0}}p\mathcal{Q}_{1}^{0}+p\mathcal{Q}_{1}^{1}+(1-p-e^{\varepsilon_{0}}p)\mathcal{Q}_{1},$
$\mathcal{R}(x_{1}^{1})$ $=p\mathcal{Q}_{1}^{0}+e^{\varepsilon_{0}}p\mathcal{Q}_{1}^{1}+(1-p-e^{\varepsilon_{0}}p)\mathcal{Q}_{1},$
@@
This new decomposition proposed in the corrected version introduces additional challenges. First, for specific randomizers, computing a tight value of $q$ is nontrivial. Second, the monotonicity of $D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p,q)\parallel P_{1}(\varepsilon_{0},p,q)\big{)}$ with respect to $p$ is not known. Consequently, we are unable to derive the desired conclusion—namely, a general upper bound applicable to any $\varepsilon_{0}$-DP local randomizer, as stated in (1). For the same reason, it remains unclear whether this decomposition necessarily yields tighter bounds than the standard clone decomposition. Moreover, the new bound $D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p,q)\parallel P_{1}(\varepsilon_{0},p,q)\big{)}$ for a specific local randomizer lacks an efficient algorithm to compute.
4 General Clone and the Optimal Bounds
@@
The hierarchy of the bounds provided by the decomposition-based methods is shown in Fig. 1.
Figure 1: Hierarchy among decomposition-based methods
Since the primary optimal decomposition derived from the blanket framework is generally highly redundant, we propose a general method for computing a simplified optimal decomposition. Furthermore, we observe that the optimal decomposition satisfies desirable compositional properties under both joint composition and parallel composition. Specifically, the optimal decomposition of a joint composition corresponds to the Cartesian product of the optimal decompositions of the sub-local randomizers, while that of a parallel composition corresponds to their weighted combination.
4.1 Definition of general clone
Definition 7 (Decomposition in the General Clone).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a local randomizer. The general clone paradigm defines a decomposition of $\mathcal{R}$ as follows:
@@ 4.1 Definition of general clone
$\forall i\in[2,n],\quad\mathcal{R}(x_{i})$ $=\sum_{j=1}^{k}c_{j}Q_{1}^{j}+\beta Q_{i},$ (2)
where $Q_{1}^{j}$ for $j=1,\dots,k$ and $Q_{i}$ for $i=2,\dots,n$ are probability distributions over $\mathbb{Y}$, and $a_{j},b_{j},c_{j},\beta$ are non-negative coefficients and satisfy:
$\sum_{j=1}^{k}a_{j}=1,\quad\sum_{j=1}^{k}b_{j}=1,\quad\beta+\sum_{j=1}^{k}c_{j}=1.$
Remark 1.
Although the index $i$ ranges over $[2,n]$, each $Q_{i}$ depends only on the corresponding input $x_{i}$. Specifically, it is defined as the “left-over” distribution:
$\text{\emph{LO}}(x_{i}):=\frac{1}{\beta}\left(\mathcal{R}(x_{i})-\sum_{j=1}^{k}c_{j}Q_{1}^{j}\right).$
For notational simplicity, we use $Q_{i}$ to denote $\text{\emph{LO}}(x_{i})$ throughout the analysis.
Remark 2.
The general clone is defined with respect to three parameters: the local randomizer $\mathcal{R}$ and two inputs $x_{1}^{0},x_{1}^{1}$. When $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ satisfies $\varepsilon_{0}$-differential privacy, a decomposition of the form given by the standard clone exists for any pair $x_{1}^{0},x_{1}^{1}\in\mathbb{X}$. However, in the general clone framework, the feasible decomposition may vary depending on the specific choice of $x_{1}^{0}$ and $x_{1}^{1}$.
The general clone paradigm characterizes the general form of decompositions used in privacy amplification analysis. It directly subsumes the decomposition used in the standard clone framework. Although the decomposition defined by the privacy blanket appears structurally different, we show that it naturally corresponds to a valid decomposition under the general clone paradigm in Section 4.2.
@@
$\Pr[A_{0}=j]=\begin{cases}a_{j}&\text{for }j\leq k,\\ 0&\text{for }j=k+1.\end{cases},\quad\Pr[A_{1}=j]=\begin{cases}b_{j}&\text{for }j\leq k,\\ 0&\text{for }j=k+1.\end{cases}$
@@
The reduction pair $(P_{0}^{\mathsf{GC}},P_{1}^{\mathsf{GC}})$ (with “$\mathsf{GC}$” denoting “General Clone”) is defined as the distributions of the histograms over $\{1,\dots,k+1\}$ generated by sampling:
  • One sample from $A_{0}$ for $P_{0}^{\mathsf{GC}}$ (or $A_{1}$ for $P_{1}^{\mathsf{GC}}$);
@@
Theorem 4 (General Clone Reduction).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a $\varepsilon_{0}$-DP local randomizer and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. For $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$, we have
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}^{\mathsf{GC}}\parallel P_{1}^{\mathsf{GC}}\big{)}.$
Proof.
The corresponding post-processing function $f^{\mathsf{GC}}$ is shown in Algorithm 3. It operates by iterating over the input. When encountering indices $1,2,\dots,k$, it samples once from each corresponding distribution $Q_{1}^{1},Q_{1}^{2},\dots,Q_{1}^{k}$. When encountering index $k+1$, it uniformly selects an element $j\in\{2,3,\dots,n\}$ that has not yet been selected, and samples once from $Q_{j}$, which is the left-over distribution corresponding to $x_{j}$. Finally, it returns all the sampled values as output.
We now prove that $f^{\mathsf{GC}}(P_{0}^{\mathsf{GC}})=\mathcal{M}_{S}(X^{0})$; the proof for $f^{\mathsf{GC}}(P_{1}^{\mathsf{GC}})=\mathcal{M}_{S}(X^{1})$ is analogous.
Recall that $P_{0}^{\mathsf{GC}}$ is a histogram over a mixture of one sample from $A_{0}$ and $n-1$ samples from $A_{2}$, while $\mathcal{M}_{S}(X^{0})$ is the shuffled mixture of $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{i})$ for $i=2,3,\dots,n$. A well-known property is that producing a histogram over outputs is equivalent to applying a random shuffling to the output multiset—each can be generated from the other [CSWS21]. For clarity of exposition, we adopt the latter interpretation.
Let $\mathcal{P}_{n}$ denote the set of all permutations over $[n]$. In $P_{0}^{\mathsf{GC}}$, the sampling is followed by a uniformly random permutation $\sigma_{1}\leftarrow_{\$}\mathcal{P}_{n}$, such that the $\sigma_{1}(1)$-th element $z_{\sigma_{1}(1)}$ is drawn from $A_{0}$. The remaining positions $\sigma_{1}(i)$, for $i=2,3,\dots,n$, are drawn from $A_{2}$. It can be justified that $f^{\mathsf{GC}}(z_{\sigma_{1}(1)})\sim\mathcal{R}(x_{1}^{0})$:
$\Pr[f^{\mathsf{GC}}(z_{\sigma_{1}(1)})=y]=\Pr[f^{\mathsf{GC}}(A_{0})=y]$ $=\sum_{m=1}^{k}\Pr[A_{0}=m]\cdot\Pr[f^{\mathsf{GC}}(m)=y]$
$=\sum_{j=1}^{k}a_{j}Q_{1}^{j}(y)=\mathcal{R}(x_{1}^{0})(y).$
An important observation is that the support of $A_{0}$ does not include the symbol $k+1$. As a result, the on-the-fly processing of $k+1$ in $f^{\mathsf{GC}}$ is equivalent to introducing a new permutation $\sigma_{2}\leftarrow_{\$}\{\sigma\in\mathcal{P}_{n}\mid\sigma(\sigma_{1}(1))=\sigma_{1}(1)\}$. Under this view, if $z_{i}=k+1$, then sampling from $Q_{\sigma_{2}(\sigma_{1}(i))}$ is equivalent to the original procedure.
We can then show that the output of $f^{\mathsf{GC}}$ at round $i$ (denoted by $f_{i}^{\mathsf{GC}}(z_{i})$ ) satisfies $f_{i}^{\mathsf{GC}}(z_{i})\sim\mathcal{R}(x_{\sigma_{2}(\sigma_{1}(i))})$, for all $i\neq\sigma_{1}(1)$:
$\Pr[f^{\mathsf{GC}}_{i}(z_{i})=y]$ $=\Pr[f^{\mathsf{GC}}_{i}(A_{2})=y]$
$=\sum_{m=1}^{k}\Pr[A_{2}=m]\cdot\Pr[f^{\mathsf{GC}}_{i}(m)=y]+\Pr[A_{2}=k+1]\cdot\Pr[f_{i}^{GC}(k+1)=y]$
$=\sum_{m=1}^{k}c_{j}Q_{1}^{m}(y)+\beta Q_{\sigma_{2}(\sigma_{1}(i))}(y)$
$=\mathcal{R}(x_{\sigma_{2}(\sigma_{1}(i))}).$
It can be shown that the composition $\sigma_{2}\circ\sigma_{1}$ is also uniformly distributed over $\mathcal{P}_{n}$. In summary, the output of $f^{\mathsf{GC}}(P_{0}^{\mathsf{GC}})$ is a random shuffling over the multiset consisting of $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{i})$ for $i=2,3,\dots,n$, which completes the proof. ∎
Algorithm 3 Post-processing function of general clone , $f^{\mathsf{GC}}$
$x_{1}^{0},x_{1}^{1},x_{2},\dots,x_{n}$; $z\in\{1,2,\dots,k+1\}^{n}$
$J\leftarrow\emptyset$
for $i=1,\dots,n$ do
     if $z_{i}=k+1$ then
         Let $j$ be a randomly and uniformly chosen element of $[2:n]\setminus J$
         $J\leftarrow J\cup\{j\}$
     end if
     Sample $y_{i}$ from
      $\begin{cases}Q_{1}^{z_{i}}&\text{if }z_{i}\in\{1,2,\dots,k\};\\ Q_{j}&\text{if }z_{i}=k+1.\end{cases}$triangleright Sample from the left-over distribution of $x_{j}$ when $z_{i}=k+1$
end for
return $y_{1},\dots,y_{n}$
4.2 Blanket is “in” the general clone
Algorithm 4 Bijection from privacy blanket to primary optimal decomposition in general clone, $f_{2}$
$(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})$, $n$, $\mathbb{Y}=\{y_{1},y_{2},\dots,y_{|\mathbb{Y}|}\}$
$H\leftarrow\{\}$triangleright $H$ is a histogram
for $y\in Y_{|\boldsymbol{b}|}$ do
     $j\leftarrow\text{index}(y)\text{ such that }y_{j}=y$
     $H[j]\leftarrow H[j]+1$triangleright Map $y=y_{j}$ in Blanket to index $j$ in GC
end for
$H[|\mathbb{Y}|+1]\leftarrow n-1-|\boldsymbol{b}|$triangleright Map $y=\bot$ in Blanket to index $|\mathbb{Y}|+1$ in GC
return $H$
Theorem 5.
For any local randomizer $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ and inputs $x_{1}^{0},x_{1}^{1}\in\mathbb{X}$, the following decomposition in the general clone paradigm is equivalent to the privacy blanket framework:
$\mathcal{R}(x_{1}^{0})$ $=\sum_{j=1}^{|\mathbb{Y}|}a_{j}\mathbbm{1}_{y_{j}},$
$\mathcal{R}(x_{1}^{1})$ $=\sum_{j=1}^{|\mathbb{Y}|}b_{j}\mathbbm{1}_{y_{j}},$
$\forall i\in[2,n],\quad\mathcal{R}(x_{i})$ $=\sum_{j=1}^{|\mathbb{Y}|}c_{j}\mathbbm{1}_{y_{j}}+\beta Q_{i},$ (3)
We refer to this as the primary optimal decomposition with respect to $\mathcal{R}$ and $(x_{1}^{0},x_{1}^{1})$.
Proof.
We construct a post-processing function $f_{2}$ (shown in Algorithm 4). The function $f_{2}$ initializes an empty set $H$, then iterates over its input. Upon encountering an element $y_{j}$, it adds the index $j$ to the set $H$. Finally, it appends $n-1-|\boldsymbol{b}|$ copies of $|\mathbb{Y}|+1$ to $H$.
We now prove that $f_{2}(P^{\mathsf{B}}_{0})=P^{\mathsf{GC}}_{0}$. The input of User 1 is sampled from $\mathcal{R}(x_{1}^{0})$ and always appears in the output of $P^{\mathsf{B}}_{0}$. By the definition of $f_{2}$, we have:
$f_{2}(\mathcal{R}(x_{1}^{0}))\sim A_{0},$
where $A_{0}$ is defined in Definition 8.
For the remaining users, each appears in the output of $P^{\mathsf{B}}_{0}$ with probability $\gamma$, and is omitted with probability $1-\gamma$. In the latter case, $f_{2}$ outputs the special index $|\mathbb{Y}|+1$. In the former case, the distribution of $f_{2}(\omega)$ matches the conditional distribution of $A_{2}$ given that $A_{2}\neq|\mathbb{Y}|+1$:
$\Pr[f_{2}(\omega)=y_{j}]=\inf_{x\in\mathbb{X}}\mathcal{R}(x)(y_{j})/\gamma=c_{j}/\sum_{m=1}^{|\mathbb{Y}|}c_{m}.$
Together, this shows that applying $P^{B}_{0}$ followed by $f_{2}$ for a remaining user is equivalent to directly sampling a random variable from $A_{2}$. Therefore, the output distribution satisfies:
$f_{2}(P^{\mathsf{B}}_{0})=P^{\mathsf{GC}}_{0}.$
The proof for $f_{2}(P^{\mathsf{B}}_{1})=P^{\mathsf{GC}}_{1}$ follows by a similar argument.
As $f_{2}$ is a bijection, a post-processing function $f_{2}^{-1}$ exists such that $f_{2}^{-1}(P^{\mathsf{GC}}_{0})=P^{\mathsf{B}}_{0},f_{2}^{-1}(P^{\mathsf{GC}}_{1})=P^{\mathsf{B}}_{1}$, thereby proving the equivalence between the two constructions. ∎
4.3 General clone is not stronger than blanket